OpenStudy (rainbow_dash):

solve the nonlinear system of equations: x^2+y^2=63 x^2-3y^2=27

OpenStudy (helder_edwin):

u could multiply the 1st equation by 3 and then add both equations.

OpenStudy (lgbasallote):

or just subtract the two equations...

OpenStudy (lgbasallote):

x^2 + y^2 = 63 - x^2 - 3y^2 = 27 ========== 0x^2 + 4y^2 = 36 4y^2 = 36 ---- --- 4 4 y^2 = 9 does that help @Rainbow_Dash ?

OpenStudy (rainbow_dash):

it did a ton, thanks!

OpenStudy (lgbasallote):

welcome ^_^